Opened 9 years ago

Closed 3 years ago

#322 closed flight processing (fixed)

GB09/04, flight day 062/2010, Coed-y-Brenin

Reported by: benj Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: iopa@… Other processors: iopa

Description (last modified by iopa)

Data location: ~arsf/arsf_data/2010/flight_data/uk/GB09_04-2010_062_Coed_Y_Brenin

Data arrived from ARSF via network transfer on 19th March 2010.

Scientific objective: Mapping Rhododendron sites using combination hyperspectral and LiDAR.

Priority: a4l

PI: Sarah Taylor

Sensors:

  • Eagle (03/09/2010 & 18/11/2010)
  • Hawk (09/12/2010 & 20/12/2010)
  • Leica LIDAR (03/09/2010)
  • RCD Digital Photography (03/09/2010)

Attachments (1)

syncOffsets (437 bytes) - added by iopa 8 years ago.
Eagle/Hawk SCT offsets

Download all attachments as: .zip

Change History (40)

comment:1 Changed 9 years ago by benj

There are 24 eagle and hawk lines but only 22 recorded on the logsheet, presumably two were re-flown. Have asked ARSF-ops which two are affected (though looking at file sizes I think the bad ones are probably 1 and 13 for Eagle and 1 and 2 for Hawk).

comment:2 Changed 9 years ago by benj

  • Milestone set to 2010 data processing completion

comment:3 Changed 9 years ago by mggr

When the photos are processed, please let Mike know - Bill is interested in copies of high-relief photos purely for internal testing purposes.

comment:4 Changed 9 years ago by pmlvis10

IPAS navigation for this project has been processed.

comment:5 Changed 9 years ago by ella

Looks like applanix navigation processing has been done. Copied to arsf workspace in Samba.

comment:6 Changed 9 years ago by iopa

  • Cc iopa@… added
  • Other processors set to iopa

Processing Hyperspectral.

comment:7 Changed 9 years ago by iopa

Hyperspectral data has been processed. PI will now be contacted to determine whether OS vector verification is required.

comment:8 Changed 9 years ago by iopa

Regarding my message above, no contact with PI is required since we already have the vectors.

comment:9 Changed 9 years ago by iopa

Eagle/Hawk SCT Offsets attached. Hawk ones have been significantly over the 0.1 positive limit in this project and above 1 second. There were also two peaks of 28.75 and 43.3 seconds.

comment:10 Changed 9 years ago by jb362

Starting delivery check

comment:11 Changed 9 years ago by jb362

Hyperspectral delivery check complete. Everything fine, although worth noting:

1) Only 60 bands for eagle data although this is because of 8,1 binning as shown in eagle header file.

2) Pixel overflow for eagle flight lines:

e2: bands 0-14
e10: bands 0-25
e11: bands 0-29
e12: bands 0-24
e13: bands 0-16
e20: bands 0-23
e22: bands 0-32

3) Hawk band 160 blank for all flight lines

comment:12 Changed 9 years ago by jb362

4) Also, you need to put the DEM in the delivery directory

comment:13 Changed 9 years ago by iopa

Hyperspectral is ready to deliver.

comment:14 Changed 9 years ago by iopa

Taking up LiDAR.

comment:15 Changed 9 years ago by iopa

  • Description modified (diff)

Taking camera processing.

comment:16 Changed 9 years ago by jb362

Delivery checking camera.

comment:17 Changed 9 years ago by jb362

Camera delivery looks fine.

comment:18 Changed 9 years ago by iopa

LiDAR processing complete and ready for checking.

comment:19 Changed 9 years ago by ella

Starting LiDAR checking.

comment:20 Changed 9 years ago by ella

Aborting delivery check.

comment:21 Changed 9 years ago by iopa

Roll offset used for LiDAR processing: 0.00089950 (rad)

comment:22 Changed 9 years ago by emca

Delivery checked LiDAR delivery at
~airborne/workspace/GB09_04-2010_062_Coed_Y_Brenin/delivery/20100901/GB09_04

Need to update logsheet to typed version and ensure the name is in the correct format. Also Julian day on the logsheet says 063 - should this be changed to 062?

Remove hidden files from delivery directory (.nfs*)


comment:23 Changed 9 years ago by iopa

  • Description modified (diff)

Dispatched USB Disk on 03/09/2010 containing Eagle, LiDAR and Photography data.

comment:24 Changed 9 years ago by anch

Rechecked after discovering that incorrect images were dispatched to PI (Due to has_sync being set to false by accident).

Everything fine with images after reprocessing with has_sync set to true.

Logsheet is still handwritten - this should be updated to a typed version

comment:25 Changed 9 years ago by iopa

  • Description modified (diff)

During email correspondence with the PI and collaborators, ARSF discovered the delivered level 1 data was not adjusted for the timing offsets and the eagle dataset was re-processed with the timing offsets for eagle as listed in the attachment syncOffsets.sync.

Logsheet for the new delivery was converted from abw to pdf and the corrected Eagle data has been dispatched to PI on 18/11/2010.

comment:26 Changed 8 years ago by iopa

Hawk is ready for checking.

comment:27 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Beginning delivery check.

comment:28 Changed 8 years ago by knpa

  • Owner changed from knpa to iopa

Delivery check complete.

All looks good and ready to deliver.

There is one thing that might be worth investigating: On band 0 for every line the rightmost samples are blacked out.

comment:29 Changed 8 years ago by benj

I noticed that on another dataset - I suspect it's a bad pixel (or more than one).

comment:30 Changed 8 years ago by iopa

  • Description modified (diff)

Hawk data dispatched on 09/12/2010.

Changed 8 years ago by iopa

Eagle/Hawk SCT offsets

comment:31 Changed 8 years ago by iopa

Hawk data is being re-processed for new calibration wavelength set.

comment:32 Changed 8 years ago by knpa

Hawk data reprocessed - beginning delivery check.

comment:33 Changed 8 years ago by knpa

Checked level 1's - ready to deliver.

comment:34 Changed 8 years ago by iopa

  • Description modified (diff)

Hawk data has been re-dispatched on 20/12/2010.

comment:35 Changed 8 years ago by iopa

Project R-Synced and workspace version has been deleted.

comment:36 Changed 7 years ago by adbe

Hyperspectral reprocessing delivery check: No problems, ready to deliver.

comment:37 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from iopa to knpa

Beginning archiving.

comment:38 Changed 3 years ago by dac

Archiving

Checked project.
Moved deliveries to 'deliveries' folder and renamed according to new convention. Previously processed hyperspectral files moved to 'previous_deliveries'

Started uploading to NEODC.

comment:39 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Processed and raw data now available from NEODC: http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB09_04/GB09_04-2010_062_Coed_Y_Brenin

Marking as closed.

Note: See TracTickets for help on using tickets.